OpenStudy (anonymous):

Identify the solution(s) of the system of equations, if any. 3x+2y=4 5x-4y=3

OpenStudy (anonymous):

First, you isolate for y in both equations. Then make the two equations equal to each other and solve for x.

OpenStudy (anonymous):

3x+2y=4 2y = 4-3x y=(4-3x)/2 And 5x-4y=3 -4y = 3-5x y=(3-5x)/-4 Put both equations together. (4-3x)/2 = (3-5x)/-4 And isolate for x. Try that out. (: